# Gyroscopically stabilized quantum spin rotors

### v1

@@ Appendix D Barnett anti-alignment
implying that $20\leq{\rm var}(\cos\gamma)\leq 1/$ 2.
Appendix E Stabilization of spin superpositions
Figure 5: (a) Occupation probability of the $|S_{1}=0\rangle$ spin state calculated from the exact dynamics given by Eq. (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) compared to the approximated solution of Eq. (43) (light blue, solid). (b) Comparison of the $\gamma$-dynamics obtained from the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) and the magnetic spin Hamiltonian (10) (light blue, solid). The nanodiamond has an ellipsoidal shape with semi-axes $l_{3}=200$ nm and $l_{1}=l_{2}=0.4l_{3}$, $B=-0.5\,\mathrm{mT}$, $\omega/2\pi=0.1\,\mathrm{MHz}$ and $\omega_{0}=\hbar/I_{\mathrm{eff}}$. The spin is initially in $\left|\sigma_{x}=-1\right>$ located at $\gamma=\pi/2$. We simulated the dynamics in the angular momentum basis with dimension $d_{m}=1729$ and rescaled $\delta$ and $g$ by a factor $10^{-8}$ due to computational constraints.
@@ Appendix E Stabilization of spin superpositions
$-\dfrac{\hbar g}{\sqrt{2}}\sin\gamma\left(\left|\chi_{+1}\right>+\left|\chi_{-1}\right>\right),$ (42a)
@@
For moderate $B$ fields, $|\omega-\gamma_{0}B|\lesssim D_{\rm nv}$, the non-magnetic state $|S_{1}=0\rangle$ remains far detuned from the magnetic states $|S_{1}=\pm\hbar\rangle$ for all orientations $\gamma$. This implies that $i\hbar\partial_{t}\left|\chi_{0}\right>\approx 0$ for all times, so that
$\left|\chi_{0}\right>\approx-\dfrac{\delta}{\sqrt{2}g}\sin\gamma\left(\left|\chi_{+1}\right>+\left|\chi_{-1}\right>\right),$ (43)
@@
Appendix F Spin-controlled rotational interference
[…]
@@ Appendix F Spin-controlled rotational interference
where we shifted the system by $-\hbar\Delta/2$ and defined
@@
$\mathbf{m}=\left[\dfrac{\Delta}{2}+\dfrac{(g+\tilde{\delta})}{4}\right]\mathbf{e}_{z}-\dfrac{g}{\sqrt{2}}\gamma\mathbf{e}_{x}$ (52)
@@
$H_{\mathrm{disp}}\simeq$ $\dfrac{p^{2}_{\gamma}}{2I_{\mathrm{eff}}}+\dfrac{\hbar(g-\tilde{\delta})}{4}\gamma^{2}$
@@
[…]
@@
[…]
@@
$I_{\gamma}$ $=\mathrm{Re}\left[\langle n_{0}|U_{\downarrow}^{\dagger}U_{\uparrow}^{\dagger}U_{\downarrow}U_{\uparrow}|n_{0}\rangle\right]$
@@
$=\mathrm{Re}\left[\left(1+\dfrac{(\Delta+4g)^{2}(1-\mathrm{e}^{2i\tau\omega_{\gamma}})}{8g(\Delta+2g)}\sinh^{2}\zeta\right)^{-\frac{1}{2}}\right].$ (57)
@@
For the second equality the Baker-Campbell-Hausdorff formula of SU(1,1) [66] was used to rewrite the evolution operators (55) [12]. For finite temperatures the expression for $I_{\gamma}$ becomes more complicated and can be found in the Supplement of Ref. [12].

### v2

@@ Appendix E Barnett anti-alignment
implying that $0\leq{\rm var}(\cos\gamma)\leq 1/2$.
Appendix F Stabilization of spin superpositions
Figure 8: (a) Occupation probability of the $|S_{1}=0\rangle$ spin state calculated from the exact dynamics given by Eq. (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) compared to the approximated solution of Eq. (51) (light blue, solid). (b) Comparison of the $\gamma$-dynamics obtained from the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) and the magnetic spin Hamiltonian (10) (light blue, solid). The nanodiamond has an ellipsoidal shape with semiaxes $l_{3}=200$ nm and $l_{1}=l_{2}=0.4l_{3}$, $B=-0.5\,\mathrm{mT}$, and $\omega/2\pi=1\,\mathrm{MHz}$. The spin is initially in $\left|\sigma_{x}=-1\right$> located at $\gamma=\pi/2$. We simulated the dynamics in the angular momentum basis with dimension $d_{m}=1729$ and rescaled $\delta$ and $g$ by a factor $10^{-8}$ due to computational constraints.
@@ Appendix F Stabilization of spin superpositions
[…]
@@
For moderate $B$ fields, $|\omega-\gamma_{0}B|\lesssim D_{\rm nv}$, the nonmagnetic state $|S_{1}=0\rangle$ remains far detuned from the magnetic states $|S_{1}=\pm\hbar\rangle$ for all orientations $\gamma$. This implies that $i\hbar\partial_{t}\left|\chi_{0}\right>\approx 0$ for all times, so that
[…]
@@
For $l_{3}=200\,\mathrm{nm}$, $l_{1}=l_{2}=0.3l_{3}$, $\omega/2\pi=1\,\mathrm{MHz}$, and $B=-100\,\mathrm{mT}$, the above expression evaluates to $\sim 5\times 10^{-4}$, justifying the perturbative treatment.
Appendix G Spin-controlled rotational interference
Starting from Eq. (Gyroscopically stabilized quantum spin rotors) we derive the dispersive Hamiltonian given in Eq. (12) by restricting the spin to the subspace $\{\left|\uparrow\right>,\left|\downarrow\right>\}=\{\left|S_{1}=+\hbar\right>,\left|S_{1}=-\hbar\right$>\}. Proceeding analogously as in Sec. F, we first harmonically approximate the system around the potential minimum $\gamma=0$ and subsequently adiabatically eliminate the $\left|\chi_{+1}\right$> state of Eq. (49). By considering magnetic field strengths $g\simeq D_{\mathrm{nv}}$, it holds $i\hbar\partial_{t}\left|\chi_{+1}\right>\approx 0$ for all times, such that
@@ Appendix G Spin-controlled rotational interference
where we shifted the system by $\hbar\Delta/2$ and defined
@@
$\mathbf{m}=\left[\dfrac{\Delta}{2}+\dfrac{(g+\tilde{\delta})}{4}\gamma^{2}\right]\mathbf{e}_{z}-\dfrac{g}{\sqrt{2}}\gamma\mathbf{e}_{x}$ (60)
@@
$H_{\mathrm{d}}\simeq$ $\dfrac{p^{2}_{\gamma}}{2I_{\mathrm{eff}}}+\dfrac{\hbar(g-\tilde{\delta})}{4}\gamma^{2}$
@@
$\dfrac{2g^{2}}{\Delta^{2}}\left<\gamma^{2}\right>\ll 1,\;\dfrac{\tilde{\delta}+g}{\Delta}\left<\gamma^{2}\right>\ll 1,$ (62)
@@
The probability of measuring spin $\left|\uparrow\!(\downarrow)\right$> after the interferometer sequence described in the main text, can be calculated as
@@
$I_{\gamma}$ $=2\mathrm{Re}\left[\langle n_{0}|U_{\downarrow}^{\dagger}U_{\uparrow}^{\dagger}U_{\downarrow}U_{\uparrow}|n_{0}\rangle\right]$
@@
$=2\mathrm{Re}\left[\left(1+\dfrac{(\Delta+4g)^{2}(1-\mathrm{e}^{2i\tau\omega_{\gamma}})}{8g(\Delta+2g)}\sinh^{2}\zeta\right)^{-\frac{1}{2}}\right],$ (65)
@@
where we used $\Delta\ll g$ and introduced $\zeta=\sqrt{\hbar}g\tau/\sqrt{I_{\rm eff}\Delta}$. For the second equality the Baker-Campbell-Hausdorff formula of SU(1,1) [81] was used to rewrite the evolution operators (63) [12]. For finite temperatures the expression for $I_{\gamma}$ becomes more complicated and can be found in the Supplement of Ref. [12].
Appendix H Rotation-induced microwave-drive detuning
Transforming the microwave drive from the lab frame into the rotor’s body frame introduces the rotational Doppler shift $\omega$ of the spin transition. In the rotating frame, the Barnett term shifts the spin splitting by $\omega$ (in our sign convention), which precisely cancels the Doppler shift of the drive, leaving the microwave $\pi/2$ pulses resonant as the particle spins [66].
If the Paul trap is slightly asymmetric, $A_{1}=-1/2-\epsilon$, $A_{2}=-1/2+\epsilon$ with $\epsilon\ll 1$, the quadrupole potential acquires a weak $\alpha$-dependence. Linearizing around $\beta=\pi/2$ ($\xi=\pi/2-\beta$), valid under gyroscopic stabilization, yields the effective potential
$V(\alpha,\beta)\approx$ $\dfrac{U_{\mathrm{ac}}^{2}(Q_{1}-Q3)^{2}}{36\omega_{\mathrm{ac}}^{2}d_{0}^{4}I}\times$
$\times\left[\dfrac{9}{4}\xi^{2}+3\epsilon\xi\cos(2\alpha)+\epsilon^{2}\sin^{2}(2\alpha)\right].$ (66)
The resulting first-order change of $p_{\alpha}$ is
$\dot{p}_{\alpha}^{(1)}=-\dfrac{U_{\mathrm{ac}}^{2}(Q_{1}-Q3)^{2}}{18\omega_{\mathrm{ac}}^{2}d_{0}^{4}I}\left[\epsilon^{2}\sin(4\alpha)-3\epsilon\xi\sin(2\alpha)\right].$ (67)
Taking $\alpha\simeq\omega t$ at zeroth order and integrating over an integer number of full rotations, this correction averages to zero. Thus, the mean rotation frequency $\omega=p_{\alpha}/I$ remains unchanged to first order. The residual instantaneous modulation of the rotation frequency $\Delta\omega(t)$ scales as $\mathcal{O}(\epsilon^{2})$ and $\mathcal{O}(\epsilon\xi)$, and is negligible for small asymmetries and small gyroscopic fluctuations ($\xi\ll 1$) on timescales of the protocol.
Appendix I Decoherence channels
In this section we quantify the dominant decoherence and heating mechanisms mentioned in the main text in more detail. All estimates are evaluated in the parameter regime of interest (rotation frequency $\omega/2\pi=1\,\mathrm{MHz}$ and interferometer duration $\tau\leq 10–30\,\mu\mathrm{s}$). We show that each decoherence channel is negligible on these timescales.
I.1 Magnetic-field noise
Since the $\gamma$-rotation is not directly driven by the RF quadrupole, the dominant decoherence channel arises from magnetic dephasing of the NV center. Experiments on isolated NV centers in nanodiamonds report coherence times of otder $T_{2}\sim 10\,\mu\mathrm{s}$ at room temperature [69, 70, 71]. Thus $\gamma$-decoherence is primarily induced by fluctuations of the homogeneous magnetic field.
The relevant noise source is Johnson-Nyquist current noise in the trap electrodes, which induces magnetic-field fluctuations that can be related to the electrode geometry via the Biot-Savart law. We model fluctuations along $\mathbf{e}_{z}$ as $\delta\mathbf{B}(t)=\delta B(t)\mathbf{e}_{z}$ with Gaussian white noise $\left<\delta B(t)\delta B(t^{\prime})\right>=A_{\mathrm{fl}}^{2}\delta(t-t^{\prime})$, where $A_{\mathrm{fl}}$ denotes the magnetic-field noise amplitude. To quantify the residual effect on $\gamma$, we describe the NV as an orientation-dependent magnetic dipole $\mathbf{m}(\Omega)=-\gamma_{0}\mathbf{S}$. For $\beta=\pi/2$, the fluctuations enter through the stochastic Hamiltonian
$H_{\mathrm{fl}}(t)=-\mathbf{m}(\Omega)\cdot\delta\mathbf{B}(t)=\hbar\gamma_{0}\delta B(t)\cos\gamma.$ (68)
Expanding the interaction-picture propagator to second order in $H_{\mathrm{fl}}(t)$ and taking the ensemble average, yields the Markov master equation
$\mathcal{L}\rho=\Gamma_{B}\left(\cos\gamma\rho\cos\gamma-\dfrac{1}{2}\{\cos^{2}\gamma,\rho\}\right),$ (69)
with $\Gamma_{B}=(\gamma_{0}A_{\mathrm{fl}})^{2}$. To estimate the induced orientational decoherence, we evaluate $\left<\gamma\right|\mathcal{L}\rho\left|\gamma^{\prime}\right>=(\gamma_{0}A_{\mathrm{fl}})^{2}(\cos\gamma-\cos\gamma^{\prime})^{2}/2$. For magnetic-field noise amplitudes of order $1\,\mathrm{nT}/\sqrt{Hz}$ [82], and for angular separations of a few zero-point amplitudes $\gamma_{\mathrm{zp}}\sim(0.5-1)\times 10^{-3}$, the resulting rate is $\lesssim 1\,\mathrm{Hz}$, well below the NV’s intrinsic dephasing rate.
I.2 Gas-collisions induced rotational decoherence
Residual gas collisions induce rotational damping [83, 84, 85]. The collision rate with gas particles of mass $m_{\mathrm{g}}$, temperature $T$ and pressure $P_{\mathrm{g}}$, can be approximated for spheroidal particles as [86, 87]
$\Gamma_{\mathrm{coll}}=l_{1}(l_{1}+l_{3})P_{\mathrm{g}}\sqrt{\dfrac{2\pi}{m_{\mathrm{g}}k_{\mathrm{B}}T}}.$ (70)
For nitrogen ($m_{\mathrm{g}}=28\,\mathrm{u}$) at room temperature and $P_{\mathrm{g}}=10^{-8}\,\mathrm{mbar}$, a nanodiamond with $l_{3}=200\,\mathrm{nm}$, $l_{1}=l_{2}=0.3l_{3}$ experiences collisions at rate $\Gamma_{\mathrm{coll}}\approx 2.8\times 10^{3}/\mathrm{s}$, corresponding to $\approx 0.03$ collisions during a $10\,\mu\mathrm{s}$ protocol. Collisional decoherence is therefore negligible and can be fully suppressed under ultra-high vacuum ($P_{\mathrm{g}}\leq 10^{-8}\,\mathrm{mbar}$).
I.3 Dipole noise and electric-field noise
Highly charged spheroidal particles tend to have negligible electric dipole moments [88] which are experimentally unobserved [48, 64]. Thus, the dominant electric noise source is quadrupole-field noise in the RF trap, known to cause motional heating [89, 90]. In our geometry, the $\gamma$-mode does not couple to the quadrupole field to leading order because the RF field primarily excites the $\beta$-mode (which is eliminated in the effective model). Consequently, quadrupole-field noise perturbs $\beta$, while leaving $\gamma$ unaffected.
Further, in a Paul trap, the RF quadrupole potential is arranged such that the homogeneous electric field vanishes at the trap center, eliminating any leading order Stark coupling to the NV spin. Small static offsets of the particle from the center, due to stray DC fields, gravity, or residual micromotion, generate a local AC field proportional to the displacement. For the sub-micron offsets typical of compensated traps, these fields are strongly reduced by diamond’s dielectric screening ($\varepsilon_{\mathrm{r}}\approx 5.7$) and remain far below the NV electric-field sensitivity ($\sim 17\,$Hz cm/V) [91, 92]. Residual Stark shifts are therefore negligible compared to both the NV’s intrinsic decoherence rate and the spin-rotation coupling scale.
I.4 Black-body radiation
[…]
$F({\gamma,\gamma^{\prime}})=\int_{0}^{\infty}\dfrac{\mathrm{d}\omega_{\mathrm{ph}}\omega_{\mathrm{ph}}^{3}}{3\pi^{2}c^{3}\varepsilon_{0}}\bar{n}(\omega_{\mathrm{ph}})\mathrm{Tr}\left[\upalpha^{\prime\prime}(\omega_{\mathrm{ph}})\left(\mathbb{1}-\mathrm{R}^{T}\mathrm{R}^{\prime}\right)\right],$ (71)
with $\mathrm{R}^{T}\mathrm{R}^{\prime}\equiv\mathrm{R}^{T}(\gamma)\mathrm{R}(\gamma^{\prime})$, $c$ the speed of light, and $\bar{n}(\omega_{\mathrm{ph}})=[\mathrm{exp}(\hbar\omega/k_{\mathrm{B}}T)-1]^{-1}$ the thermal photon occupation number. Assuming the polarizability to be frequency-independent, $\upalpha^{\prime\prime}(\omega_{\mathrm{ph}})=\alpha_{\perp}^{\prime\prime}\mathbb{1}+(\alpha_{\parallel}^{\prime\prime}-\alpha_{\perp}^{\prime\prime})\mathbf{e}_{z}\otimes\mathbf{e}_{z}$, where $\alpha_{\perp}^{\prime\prime}$, $\alpha_{\parallel}^{\prime\prime}$ are the principal values along the minor and major axes, respectively, the localization rate reduces to
$F({\gamma,\gamma^{\prime}})=\Gamma_{\mathrm{ph}}[1-\cos(\gamma-\gamma^{\prime})],$ (72)
with the photon emission rate
$\Gamma_{\mathrm{ph}}=\dfrac{2\pi^{2}\alpha_{\perp}^{\prime\prime}k_{\mathrm{B}}^{4}T^{4}}{45c^{3}\hbar^{4}\varepsilon_{0}}.$ (73)
Under ambient conditions ($T\simeq 300$ K), the emission rate is $\Gamma_{\mathrm{ph}}/2\pi\approx 7.2\,\mathrm{MHz}$ for a transverse polarizability of $\alpha^{\prime\prime}_{\perp}\simeq 1\times 10^{-32}\,\mathrm{C}\mathrm{m}^{2}/\mathrm{V}$ [93]. The relevant decoherence depends on the angular separation in the superposition, as determined by the amount of squeezing. The interferometer produces superpositions of a few zero-point amplitudes $\gamma_{\mathrm{zp}}\sim(0.5-1)\times 10^{-3}$, so that $F(\gamma,\gamma^{\prime})/2\pi\lesssim 30\,\mathrm{Hz}$, well below the interferometer timescale.
Rayleigh scattering of thermal photons does not contribute because the $\gamma$-rotation is symmetric under the angular scattering distribution [94].